# Reinforcement Learning for Durable Algorithmic Recourse

### v1

@@ 1 Introduction
Actionable recommendations based on counterfactual explanations enable individuals to improve their chances of acceptance in the future (Karimi et al., 2021; Upadhyay et al., 2025). However, shifts in the training data, prediction model, or applicant pool can render such recommendations invalid over time, leading to situations where individuals who follow the suggested changes—often at significant time, labor, or financial costs—still get rejected (Upadhyay et al., 2021; Fonseca et al., 2023).
@@
This concern has motivated the development of robust recourse methods that seek to remain effective in dynamic settings, contingent on the socio-technical context in which the system operates and responsive to the evolving conditions of the decision-making system and its environment (Upadhyay et al., 2021; Dominguez-Olmedo et al., 2022; Pawelczyk et al., 2023a). In particular, when considering limited-resource, competitive settings, it becomes essential to account for and manage the feedback effects of recourse on the applicant pool (Fonseca et al., 2023). Namely, as candidates repeatedly apply after attempting to follow the recommendations, the decision threshold may shift, potentially leading to a high rate of invalidity (Bell et al., 2024). While prior work has identified this issue and emphasized the limitations of existing recourse methods under such endogenous dynamics (Fonseca et al., 2023; Segal et al., 2024), no comprehensive solution has yet been proposed.
@@
  • We introduce a comprehensive, time-aware recourse framework that models a competitive, limited-resource setting in which recommendations are issued. Our environment captures varying feature-modification difficulties and delays between candidate reapplications, thus reflecting complex human behavior and contextual constraints.
@@ 2 Related work
Algorithmic recourse emerged in response to concerns about the opacity of automated decision-making, particularly in the context of the GDPR’s right to an explanation. A foundational contribution came from Wachter et al. (2017), who introduced counterfactual explanations as a way to help individuals understand and contest model outcomes; they casted recourse as an optimization problem, where the goal was identifying the smallest set of feature changes that would alter a decision. Building on this idea, Ustun et al. (2019) formalized recourse in terms of practical costs, proposing integer programming methods to generate actionable changes for linear classifiers. Subsequent work generalized these approaches to incorporate richer objectives (Dandl et al., 2020; Mothilal et al., 2020; Cheon et al., 2025; Rasouli and Yu, 2024; Rawal and Lakkaraju, 2020). Among these, Mothilal et al. (2020) introduced DiCE, which generates diverse sets of feasible counterfactuals.
@@
Another strand of work examines the robustness of recourse in dynamic environments (Altmeyer et al., 2023; Yang et al., 2025; Kayastha et al., 2024; Stundefinedpka et al., 2025; De Toni et al., 2025). Existing work has largely focused on exogenous model shifts in non-competitive settings (Upadhyay et al., 2021; Pawelczyk et al., 2023a; Guyomard et al., 2023; Nguyen et al., 2023). Upadhyay et al. (2021) proposed a min–max optimization framework that ensures recourse validity under worst-case perturbations to model parameters and inputs. Dominguez-Olmedo et al. (2022) introduced adversarially robust strategies for counterfactual generation, while Pawelczyk et al. (2023b) highlighted the trade-off between robustness and compliance with the right to be forgotten.
Fonseca et al. (2023) and Bell et al. (2024) explore endogenous population shifts induced by recourse in competitive environments. They introduce an agent-based simulation framework to analyze how applicant competition affects recourse validity. They conclude that the state-of-the-art approach of pushing rejected candidates towards the last-seen decision threshold is ineffective, as it leads to high values of invalidity. While these works highlight the challenge of maintaining valid recourse under competition, they stop short of offering concrete solutions.
Recent research incorporates reinforcement learning into algorithmic recourse. For instance, De Toni et al. (2024) leverage RL to learn individual preferences and generate tailored recourse plans. Going further, Kanamori et al. (2025) apply RL to the concept of improvement (König et al., 2023), ensuring recommendations not only increase the chance of acceptance but also positively affect the system where the recourse is issued. Other work highlights the role of risk and imperfect user execution. To address this, Wu et al. (2024) use RL to balance cost and risk, providing policies that allow individuals to select safer options, while Xuan et al. (2025) use RL to generate robust action trajectories that account for imperfect execution.
@@ 3.1 Simulation Environment
For candidates who remain, each modification on each feature $i$ is implemented successfully with a probability of success that depends on the change magnitude, a feature-specific difficulty parameter $d_{i}\in[0,1]$ (Lievens et al., 2005), and a global difficulty parameter $\beta$. In addition, each candidate has a reapplication probability, which increases with (i) self-confidence, measured by the extent to which recommended changes were applied, and (ii) urgency, determined by the time since the last application (Grbic and Roskovensky, 2012). Candidates may delay reapplication for up to $T$ steps, consistent with the guaranteed validity of the recommendation.
These extensions improve upon prior simulations (Fonseca et al., 2023; Bell et al., 2024), which assumed (i) zero dropout probability, meaning that candidates only left once accepted, (ii) uniform modification difficulty across features, and (iii) immediate reapplication without guarantees of recommendation duration. A detailed specification of these mechanisms is provided in Appendix B.
@@ 3.2 Reinforcement learning setting
Action $a_{t}$. The agent’s action at time $t$ is defined as: $a_{t}=X_{t}^{\text{CF}}$, where $X_{t}^{\text{CF}}$ is a matrix of counterfactual feature vectors, each corresponding to a rejected candidate. These vectors represent the feature configurations that, if adopted, would lead to acceptance, within a time window of $T$ steps. The action space $\mathcal{A}$ is continuous and of variable dimension.
@@
Observation $o_{t}$ and Observation Function $O(o_{t}\mid s_{t},a_{t-1})$. The agent has partial observability of the environment, and the observation function specifies how this partial view is derived from the true latent state $s_{t}$ and the previous action $a_{t-1}$. Formally, the observation includes only the elements of $s_{t}$ corresponding to the current applicants: $o_{t}=(X_{t}^{\text{F}},\mathcal{I}_{t}^{\text{F}}).$ The observation space $\Omega$ is continuous and can have variable dimensionality depending on the number of applicants at time $t$.
@@ 4 Reinforcement learning Solution
Directly learning the full counterfactual matrix $X_{t}^{\text{CF}}$ is computationally expensive due to its high dimensionality and variable size. To address this, we adopt a hierarchical approach that separates counterfactual generation from goal selection, explicitly modeling the dependency between the two.
@@
The environment evolves according to these recommendations, making the training of $\mu$ inherently dependent on $\phi$.
In this hierarchical setup, $\mu$ decides what score to aim for, while $\phi$ determines how to modify the features to reach that score. Pre-training $\phi$ reduces the computational complexity and stabilizes the training of $\mu$. While $\mu$ is primarily responsible for the trade-off between Recourse Reliability (Equation 2) and Recourse Feasibility (Equation 3) in the reward function, the Pareto efficiency of this trade-off largely depends on $\phi$, as the feasibility of a recommendation critically depends on the trajectory taken to reach the target score. Additionally, $\phi$ indirectly optimizes the Gini index (Equation 1), by providing recommendations that approximately lead to the same score for all candidates.
@@ 4.1 Training of the Recourse Recommender Policy
The recourse recommender policy $\phi$ is trained in a simplified environment derived from the setting introduced earlier. The state at time $t$ is defined as $s_{t}=(x_{t}^{\text{F}},g)$, where $x_{t}^{\text{F}}$ is the feature vector of a single candidate and $g$ a target score. The action is the counterfactual feature vector $a_{t}=x_{t}^{\text{CF}}$.
@@
Recommendations are evaluated on two criteria: (i) accuracy: how closely $M(x_{t}^{\text{CF}})$ approaches $g$, and (ii) cost: the effort required to modify $x_{t}^{\text{F}}$ into $x_{t}^{\text{CF}}$. The accuracy objective ensures that the recourse recommender can generate paths toward arbitrary targets, and that—once paired with the recommender—it leads to counterfactual scores that are consistent with goal scores, thereby improving the Gini-based reward (Equation 1). Formally, accuracy is measured as
@@
For optimization, we employ the Soft Actor-Critic (SAC) algorithm (Haarnoja et al., 2018), a model-free, off-policy method well-suited for continuous action spaces. In our work, $\phi$ is trained online, interacting directly with the environment; the same procedure can also be executed offline if a sufficiently rich dataset of candidate features, recommendations, and outcomes is available.
@@ 5.1 Setup
In this section, we present the performance evaluation of our method relative to established baselines from the literature. Additionally, we analyze how environmental constraints and design choices influence achievable performance. Our approach is compared against three widely used baselines for non-causal recourse (Ustun et al., 2019; Wachter et al., 2017; Mothilal et al., 2020) (hereafter called Ustun, Wachter, and DiCE). Some of these methods have also been adopted as baselines in recent studies on recourse under competition (Fonseca et al., 2023; Bell et al., 2024); in our framework, they serve as alternatives to the recourse recommender policy $\phi$.
@@
We evaluate all methods under four experimental conditions, varying the recourse horizon ($T\in\{1,5\}$) and the setting difficulty ($\beta\in\{0.05,0.01\}$). The reward coefficients $(\alpha,\tau)$, which govern the trade-off between Recourse Reliability and Recourse Feasibility, are chosen to produce Pareto frontiers spanning recourse reliability values $\text{RR}_{t}^{T}$ approximately in $(0.20,0.95)$.
Details on training and evaluation are illustrated in Appendix D. Since baselines tend to overlap, we depict only Ustun for plot readability. Analogous results, including comparisons with Wachter and DiCE, and analysis of the Gini Index of all methods, are provided in Appendix E.
5.2 Results
@@
Figure 2: Comparison of Pareto fronts of our method (blue), the hybrid method based on Ustun’s approach (orange), and the baseline using Ustun’s approach (green), across four settings with $T\in\{1,5\}$ and $\beta\in\{0.05,0.01\}$. Pareto fronts plot Recourse Reliability $\text{RR}_{t}^{T}$ and Recourse Feasibility $\text{RF}_{t}^{T}$, each averaged over ten evaluation episodes.
Figure 2 shows Pareto plots for Recourse Feasibility ($\text{RF}_{t}^{T}$) and Recourse Reliability ($\text{RR}_{t}^{T}$) across all four experimental settings, averaged over ten evaluation episodes. Each point on a Pareto front corresponds to a different trained predictor $\mu$ with varying values of the parameters $\alpha$ and $\tau$.
Impact of the time horizon $T$. Comparing the top and bottom plots in Figure 2, we observe that the value of $T$ strongly affects the validity-feasibility trade-off: achieving high validity requires policies with lower feasibility as $T$ increases. Guaranteeing recourse over a longer horizon imposes a more stringent requirement, forcing the agent to recommend more challenging feature changes.
Figure 3 further highlights this phenomenon, by plotting the average Recourse Feasibility $\text{RF}_{t}^{T}$, fixing $\text{RR}_{t}^{T}=0.95$ and $\beta=0.05$, for $T\in[1,5]$. As noticed, feasibility must decrease to guarantee large reliability over an increasing time horizon $T$.
@@
Figure 3: Recourse Feasibility $\text{RF}_{t}^{T}$, for a fixed value of Recourse Reliability $\text{RR}_{t}^{T}$ ($\approx 0.95$), and $\beta=0.05$, varying $T\in[1,5]$, for our method and the hybrid (based on Ustun’s approach).
@@
Impact of the setting difficulty $\beta$. Comparing the left and right panels of Figure 2 shows that the value of $\beta$ strongly shapes the attainable trade-off between $\text{RF}_{t}^{T}$ and $\text{RR}_{t}^{T}$.
We recall that $\beta$ scales the probability of successfully implementing feature modifications: higher values correspond to higher probabilities of success, while lower values make modifications more difficult. In both scenarios, to ensure high reliability, the agent recommends relatively high target scores, that push reapplying candidates above new applicants. For large $\beta$, this strategy has a moderate negative impact on $\text{RF}_{t}^{T}$, since even challenging modifications remain feasible. In contrast, for low $\beta$, the same strategy yields a much sharper trade-off, as many candidates are unable to realize the recommended changes.
@@
This analysis reveals another intrinsic limitation of recourse in resource-constrained environments. When the means for improvement are inherently difficult (low $\beta$), it is challenging to devise recommendations that are both likely to be implemented and sufficient to guarantee a positive outcome. Consequently, practitioners must carefully prioritize among these desiderata.
Comparison of baseline, hybrid, and our approach. Across all subplots in Figure 2, the baseline approach (green) attains low reliability ($\text{RR}_{t}^{T}\approx 0.4$) while favoring feasibility. This outcome reflects the limitations of simply using the last observed decision threshold as the target score instead of tailoring it to an evolving environment. In contrast, our predictor $\mu$ can be plugged into any recourse recommender, such as Ustun, explicitly governing this trade-off (orange). These results highlight the advantage of an RL-based predictor over a simplistic fixed threshold policy.
Figure 2 also compares our method (blue) with the hybrid approach (orange). Our approach achieves Pareto optimality across all four experimental settings. The key distinction lies in the recommendation strategy: while Ustun’s method selects recommendations based solely on minimal feature changes, our policy $\phi$ explicitly accounts for feature modification difficulties, prioritizing changes to easier features, thus resulting in more feasible recourse paths. Hybrid methods based on Wachter and DiCE achieve a performance similar to Ustun, as illustrated in Appendix E.
We further note that the relative advantage of our method depends on the setting parameters. When the complexity of the problem—encoded in the parameters $\beta$ and TT—increases, the Pareto front of our method moves closer to that of the hybrid. One possible interpretation is that, while in favorable conditions ($T=1$, $\beta=0.05$) more attainable recourse paths directly correspond to a higher portion of implementing candidates (for the same validity), in more constrained environments (e.g., $T=5$ or $\beta=0.01$) this mapping is less immediate, limiting the achievable gain.
Even under these stricter conditions, however, our method continues to provide robust improvements and maintains good trade-offs. This indicates that our approach remains effective over longer horizons $T$, highlighting its practical advantage in dynamic, multi-step settings.
6 Dicussion and Conlusion

### v2

@@ 1 Introduction
Actionable recommendations based on counterfactual explanations enable individuals to improve their chances of acceptance in the future (Karimi et al., 2021; Upadhyay et al., 2025). However, shifts in the training data, prediction model, or applicant pool can render such recommendations invalid over time, leading to situations where individuals who follow the suggested changes—often at significant time, labor, or financial costs—still get rejected (upadhyay2021towards; Fonseca et al., 2023).
@@
This concern has motivated the development of robust recourse methods that seek to remain effective in dynamic settings, contingent on the socio-technical context in which the system operates and responsive to the evolving conditions of the decision-making system and its environment (upadhyay2021towards; Dominguez-Olmedo et al., 2022; Pawelczyk et al., 2023a). In particular, when considering limited-resource, competitive settings, it becomes essential to account for and manage the feedback effects of recourse on the applicant pool (Fonseca et al., 2023). Namely, as candidates repeatedly apply after attempting to follow the recommendations, the decision threshold may shift, potentially leading to a high rate of invalidity (Bell et al., 2024); we exemplify this challenge in Appendix A with a motivating scenario. While prior work has identified this issue and emphasized the limitations of existing recourse methods under such endogenous dynamics (Fonseca et al., 2023; Segal et al., 2024), no comprehensive solution has yet been proposed.
@@
  • We introduce a comprehensive, time-aware recourse framework that models a competitive, limited-resource setting in which recommendations are issued. Our environment captures feature-modification difficulties and delays between candidate reapplications, reflecting complex human behavior and contextual constraints.
@@ 2 Related work
Foundational Works on Algorithmic Recourse. Algorithmic recourse emerged in response to concerns about the opacity of automated decision-making, particularly in the context of the GDPR’s right to an explanation. A foundational contribution came from Wachter et al. (2017), who introduced counterfactual explanations as a way to help individuals understand and contest model outcomes; they casted recourse as an optimization problem, where the goal was identifying the smallest set of feature changes that would alter a decision. Building on this idea, Ustun et al. (2019) formalized recourse in terms of practical costs, proposing integer programming methods to generate actionable changes for linear classifiers. Subsequent work generalized these approaches to incorporate richer objectives (Dandl et al., 2020; Mothilal et al., 2020; Cheon et al., 2025; Rasouli and Yu, 2024; Rawal and Lakkaraju, 2020). Among these, Mothilal et al. (2020) introduced DiCE, which generates diverse sets of feasible counterfactuals.
@@
Dynamic recourse (exogenous shifts). Another strand of work examines the robustness of recourse in dynamic environments (Altmeyer et al., 2023; Yang et al., 2025; Kayastha et al., 2024; Stundefinedpka et al., 2025; De Toni et al., 2025). Existing work has largely focused on exogenous model shifts in non-competitive settings (upadhyay2021towards; Pawelczyk et al., 2023a; Guyomard et al., 2023; Nguyen et al., 2023). upadhyay2021towards proposed a min–max optimization framework that ensures recourse validity under worst-case perturbations to model parameters and inputs. Dominguez-Olmedo et al. (2022) introduced adversarially robust strategies for counterfactual generation, while Pawelczyk et al. (2023b) highlighted the trade-off between robustness and compliance with the right to be forgotten.
Reinforcement Learning solutions. Recent research incorporates reinforcement learning into algorithmic recourse. For instance, De Toni et al. (2024) leverage RL to learn individual preferences and generate tailored recourse plans. Going further, Kanamori et al. (2025) apply RL to the concept of improvement (König et al., 2023), ensuring recommendations not only increase the chance of acceptance but also positively affect the system where the recourse is issued. Other work highlights the role of risk and imperfect user execution: Wu et al. (2024) use RL to balance cost and risk, providing policies that allow individuals to select safer options, while Xuan et al. (2025) use RL to generate robust action trajectories that account for imperfect execution.
Competitive, limited-resourse setting. Fonseca et al. (2023) and Bell et al. (2024) explore endogenous population shifts induced by recourse in competitive environments. They introduce an agent-based simulation framework to analyze how applicant competition affects recourse validity. They conclude that the state-of-the-art approach of pushing rejected candidates towards the last-seen decision threshold is ineffective, as it leads to high values of invalidity. Although these works highlight the challenge of maintaining valid recourse under competition, they stop short of offering concrete solutions.
@@ 3.1 Simulation Environment
For candidates who remain, each modification on each feature $i$ is implemented successfully with a probability of success that depends on the change magnitude, a feature-specific difficulty parameter $d_{i}\in[0,1]$ (Lievens et al., 2005), and a global difficulty parameter $\beta$. In addition, each candidate has a reapplication probability, which increases with (i) self-confidence, measured by the extent to which recommended changes were applied, and (ii) urgency, determined by the time since the last application (Grbic and Roskovensky, 2012). Candidates may delay reapplication for up to $T$ steps, consistent with the guaranteed validity of the recommendation. These extensions improve upon prior simulations (Fonseca et al., 2023; Bell et al., 2024), which assumed (i) zero dropout probability, meaning that candidates only left once accepted, (ii) uniform modification difficulty across features, and (iii) immediate reapplication without guarantees on recommendation duration. A detailed specification of the mechanisms here presented and the improvements with respect to previous work is provided in Appendix B.
@@ 3.2 Reinforcement learning setting
Action $a_{t}$. The agent’s action at time $t$ is defined as: $a_{t}=X_{t}^{\text{CF}}$, where $X_{t}^{\text{CF}}$ is a matrix of counterfactual feature vectors, each corresponding to a rejected candidate. These vectors represent the configurations that, if adopted, would lead to acceptance, within a time window of $T$ steps. The action space $\mathcal{A}$ is continuous and of variable dimension.
@@
Observation $o_{t}$ and Observation Function $O(o_{t}\mid s_{t},a_{t-1})$. The agent has partial observability of the environment, and the observation function specifies how this partial view is derived from the true state $s_{t}$ and the action $a_{t-1}$. Formally, the observation includes only the elements of $s_{t}$ corresponding to the current applicants: $o_{t}=(X_{t}^{\text{F}},\mathcal{I}_{t}^{\text{F}}).$ The observation space $\Omega$ is continuous and has variable dimension depending on the number of applicants at time $t$.
@@ 4 Reinforcement learning Solution
In the previous section, we interpreted the simulation environment through the lens of reinforcement learning, framing the recourse task as learning a policy that yields valid, durable, and feasible recommendations. In this section, we describe the strategy used to search for the optimal policy within this environment.
Directly learning the counterfactual matrix $X_{t}^{\text{CF}}$ is computationally expensive due to its high dimensionality and variable size. To address this, we adopt a hierarchical approach that separates counterfactual generation from goal selection, explicitly modeling their inter-dependency.
@@
The environment evolves according to these recommendations, making the training of $\mu$ dependent on $\phi$.
In this hierarchical setup, $\mu$ decides what score to aim for, while $\phi$ determines how to modify the features to reach that score; pre-training $\phi$ reduces the computational complexity and stabilizes the training of $\mu$. While $\mu$ is primarily responsible for the trade-off between Recourse Reliability (Equation 2) and Recourse Feasibility (Equation 3) in the reward function, the Pareto efficiency of this trade-off largely depends on $\phi$, as the feasibility of a recommendation critically depends on the trajectory taken to reach the target score. Additionally, $\phi$ indirectly optimizes the Gini index (Equation 1), by providing recommendations that approximately lead to the same score for all candidates.
@@ 4.1 Training of the Recourse Recommender Policy
The recourse recommender policy $\phi$ is trained in a simplified environment derived from the setting introduced earlier. The state at time $t$ is defined as $s_{t}=(x_{t}^{\text{F}},g)$, where $x_{t}^{\text{F}}$ is the feature vector of a single candidate and $g$ a target score. The action is the counterfactual feature vector $a_{t}=x_{t}^{\text{CF}}$. Although the deployment setting of the policy remains unchanged, with respect to the setting presented in Section 3.2, the POMDP used for training is modified to pursue a different objective, which is achieving a predefined target score, and to focus exclusively on a single candidate, which is sufficient for the intended task. This reduction significantly lowers the computational burden.
@@
Recommendations are evaluated on two criteria: (i) accuracy: how closely $M(x_{t}^{\text{CF}})$ approaches $g$, and (ii) cost: the effort required to modify $x_{t}^{\text{F}}$ into $x_{t}^{\text{CF}}$. The accuracy objective ensures that the recommender can generate paths toward arbitrary targets, and that—once paired with the predictor—it leads to counterfactual scores that are consistent with goal scores, thereby improving the Gini-based reward (Equation 1). Formally, accuracy is measured as:
@@
For optimization, we employ the Soft Actor-Critic (SAC) algorithm (Haarnoja et al., 2018), a model-free, off-policy method well-suited for continuous action spaces. In our work, $\phi$ is trained online, interacting directly with the environment; the same procedure can also be executed offline if a sufficiently rich dataset, related to another recourse system employed in this setting and containing information on how candidates respond to recommendations, is available.
@@ 5.1 Setup
In this section, we present the performance evaluation of our method relative to established baselines from the literature. Additionally, we analyze how environmental constraints and design choices influence achievable performance. Our approach is compared against three widely used baselines for non-causal recourse (Ustun et al., 2019; Wachter et al., 2017; Mothilal et al., 2020) (hereafter called Ustun, Wachter, and DiCE). Some of these methods have also been adopted as baselines in recent studies on recourse under competition (Fonseca et al., 2023; Bell et al., 2024); in our framework, they serve as alternatives to the recommender $\phi$.
@@
We further compare our approach with the method of Dominguez-Olmedo et al. (2022), referred to as Adversarially Robust Recourse (ARR). ARR accounts for adversarial perturbations to an individual’s features; under a linear model $M(\cdot)$ with independent features, the ARR objective reduces to adjusting the target score by a margin $\varepsilon$, and then deriving counterfactuals to reach this adjusted target using the same procedure as Ustun et al. (2019).
@@
Figure 2: Comparison of Pareto fronts of our method (blue), the hybrid method based on Ustun’s approach (orange), the baseline using Ustun’s approach (green), and the ARR method (red), across four settings with $T\in\{1,5\}$ and $\beta\in\{0.05,0.01\}$. Pareto fronts plot Recourse Reliability $\text{RR}_{t}^{T}$ and Recourse Feasibility $\text{RF}_{t}^{T}$, each averaged over ten evaluation episodes. The gray line at $RR_{t}=0.8$ denotes the high reliability threshold, distinguishing configurations that achieve desirable recourse reliability.
We evaluate all methods under four experimental conditions, varying the recourse horizon ($T\in\{1,5\}$) and the setting difficulty ($\beta\in\{0.05,0.01\}$). The reward coefficients $(\alpha,\tau)$, and the robustness coefficient $\varepsilon$, which govern the trade-off between Recourse Reliability and Recourse Feasibility, are chosen to produce Pareto frontiers spanning recourse reliability values $\text{RR}_{t}^{T}$ approximately in $(0.20,0.95)$.
Details on the training and evaluation procedures are provided in Appendix D. For plot readability, we depict only results for Ustun and ARR in this section; analogous results—including comparisons with Wachter and DiCE, as well as an analysis of the Gini Index for all methods—are deferred to Appendix E. Finally, Appendix G presents results from a simulation that assesses the applicability of our method in a real-world setting, in which candidates’ initial features are drawn from the German Credit dataset.111archive.ics.uci.edu/dataset/144/statlog+german+credit+data
5.2 Results
Figure 2 shows Pareto plots for Recourse Feasibility ($\text{RF}_{t}^{T}$) and Recourse Reliability ($\text{RR}_{t}^{T}$) across all four experimental settings, averaged over ten evaluation episodes. Regarding our method and the hybrid approach, each point on a Pareto front corresponds to a different trained predictor $\mu$ with varying values of the parameters $\alpha$ and $\tau$. In the case of ARR, each point corresponds to a different value of $\varepsilon$. The horizontal gray line at $RR_{t}=0.8$ indicates the high reliability threshold, highlighting simulations that achieve a desirable level of recourse reliability.
Impact of the time horizon $T$. Comparing the top and bottom plots in Figure 2, we observe that the value of $T$ strongly affects the validity-feasibility trade-off: achieving high validity requires policies with lower feasibility as $T$ increases. Guaranteeing recourse over a longer horizon imposes a more stringent requirement, forcing the agent to recommend more challenging feature changes. Figure 3 further highlights this phenomenon, by plotting the average Recourse Feasibility $\text{RF}_{t}^{T}$, fixing $\text{RR}_{t}^{T}=0.95$ and $\beta=0.05$, for $T\in[1,5]$. As noticed, feasibility must decrease to guarantee large reliability over an increasing time horizon.
@@ 5.2 Results
Figure 3: Recourse Feasibility $\text{RF}_{t}^{T}$, for a fixed value of Recourse Reliability $\text{RR}_{t}^{T}$ ($\approx 0.95$), and $\beta=0.05$, varying $T\in[1,5]$, for our method, the hybrid (based on Ustun’s approach), and ARR.
Impact of the setting difficulty $\beta$. Comparing the left and right panels of Figure 2 shows that the value of $\beta$ strongly shapes the attainable trade-off between $\text{RF}_{t}^{T}$ and $\text{RR}_{t}^{T}$. We recall that $\beta$ scales the probability of successfully implementing feature modifications: higher values correspond to higher probabilities of success, while lower values make modifications more difficult. In both scenarios, to ensure high reliability, the agent recommends relatively high target scores, that push reapplying candidates above new applicants. For large $\beta$, this strategy has a moderate negative impact on $\text{RF}_{t}^{T}$, since even challenging modifications remain feasible. In contrast, for low $\beta$, the same strategy yields a much sharper trade-off, as many candidates are unable to realize the recommended changes. This analysis reveals another intrinsic limitation of recourse in resource-constrained environments. When the means for improvement are inherently difficult (low $\beta$), it is challenging to devise recommendations that are both likely to be implemented and sufficient to guarantee a positive outcome. Consequently, practitioners must carefully prioritize among these desiderata.
@@
Comparison of baseline, hybrid, ARR, and our approach. Across all subplots in Figure 2, the baseline approach (green) attains low reliability ($\text{RR}_{t}^{T}\approx 0.4$) while favoring feasibility. This outcome reflects the limitations of simply using the last observed decision threshold as the target score instead of tailoring it to an evolving environment. In contrast, our predictor $\mu$ can be plugged into any recourse recommender, such as Ustun, explicitly governing this trade-off (orange). These results highlight the advantage of an RL-based predictor over a simplistic fixed threshold policy. Comparing ARR (red) to the baseline (green), we observe that robustifying the target score based on a fixed parameter $\varepsilon$ improves recourse reliability, at the cost of reduced feasibility. However, the hybrid method (orange) Pareto-dominates ARR in all four settings, particularly in high-reliability regimes ($RR_{t}$ above the high reliability threshold). This demonstrates that environment-aware target adjustments outperform naive approaches that ignore the candidates’ behavior. A more in-depth analysis of this is provided in Appendix F.
Figure 2 also compares our method (blue) with the hybrid approach (orange). Our approach achieves Pareto optimality across all four settings. The key distinction lies in the recommendation strategy: while Ustun selects recommendations based solely on minimal feature changes, our policy $\phi$ explicitly accounts for feature difficulties, prioritizing changes to easier features, resulting in more feasible recourse paths. Hybrid methods based on Wachter and DiCE achieve a performance similar to Ustun, as illustrated in Appendix E. Overall, these results highlight the effectiveness of our method in achieving more favorable reliability–feasibility trade-offs in dynamic, multi-step settings.
6 Discussion and Conclusions